OpenStudy (hockeychick23):

Suppose that many samples are taken from a population whose standard deviation is 5.5. If the size of each sample is 64, which of the following will most likely be the standard deviation of the standard means? A. 0.92 B. 0.69 C. 0.61 D. 0.79

OpenStudy (kropot72):

The standard deviation of the sample means is given by: \[\large \frac{\sigma}{\sqrt{n}}\] where sigma is the standard deviation of the population, and n is the sample size.

OpenStudy (hockeychick23):

ok so 5.5/8= .6875 is what i got

OpenStudy (kropot72):

Which rounds to 0.69.